OpenStudy (anonymous):

find the distance of the parallel lines: y=-2x+10 and y=-2x-5

geerky42 (geerky42):

How I would approach this problem: I will create line that is perpendicular to these lines; maybe \(y=\dfrac{1}{2}x\) Then I would find intersection point between \(y=\dfrac{1}{2}x\) and \(y = -2x+10\) And intersection point between \(y=\dfrac{1}{2}x\) and \(y = -2x-5\) Now we have two points. Find the distance between them using distance formula: \(d = \sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) Lot of works, I know. Sorry, but does that help?

OpenStudy (anonymous):

hi sorry but i still dont get it

OpenStudy (anonymous):

i already know the formula but how do you find the intersection point?

geerky42 (geerky42):

By solve for x. Just set two equations to be equal to each other.

geerky42 (geerky42):

So you know what to do?

OpenStudy (anonymous):

yeah i think so..anyway ty for your help